# Electron impact partial ionization cross sections: R-Carvone, 2-Butanol, Imidazole, and 2-Nitroimidazole

### 20 Dec 2023 v1

@@ III.3 Imidazole (IMI) and 2-Nitroimidazole (2NI)
In Figure 5(b), we have compared all the TICS from the existing literature along with our data. Our BEB TICS compares very well with the BEB TICS presented by Tejas et al.Jani et al. (2023), albeit their TICS calculated using the csp-ic method has a higher magnitude in comparison to BEB data, especially at higher energies after the peak. The magnitudes of the sums of the PICS calculated using both the MSD and the m-BEB-B methods are lower than the TICS due to the lack of information on all the fragments and the AEs as our methods are dependent on the AEs and mass spectrum data of the fragments of the parent molecule.
@@
The PICS of 2NI is shown in Figure. 6(a), which contains the PICS of 8 fragments. The smallest fragment is of m/z of 28 which has an AE of 11.44 eV and the biggest fragment is having the m/z of 97. The cation has two AEs, the fragment with the relatively lowest AE of 13.0 eV is marked with the superscript of a, and the other fragment with an AE of 16.29 eV is identified with the superscript of b, this is solely to distinguish. Like the PICS of IMI a similar trend is seen here, both the MSD PICS and the m-BEB-B PICS show a good agreement from the ionization threshold to 100 eV, after which the PICS of the MSD method falls short on comparison with the m-BEB-B PICS, but their magnitudes remain similar.
@@
The TICS of 2NI is shown in Figure. 6(b) with the BEB model is compared with the TICS obtained after summing the cross sections of all the available fragments in the MSD and m-BEB-B methods. The TICS do not match very well with the BEB TICS, as we don’t have information on the AEs of all the fragments that appeared in the EIMS. All the details of the parameters required to compute PICS are given in the Table. 4 which contains the theoretical BRs, the experimental BRs, the scaling factors, and the structural factors along with the maximum CS.
@@ IV Conclusion
In the present work, we have theoretically computed the partial and total ionization cross sections of R-carvone, 2-butanol, imidazole, and 2-nitroimidazole using the MSD and m-BEB-B methods in combination with the BEB model. We obtained a fairly good agreement of our theoretical PICS data of various fragments with the experimental PICS for R-carvone and 2-butanol. Also, the TICS which is obtained as the sum of all the PICS of the fragments in the MSD and m-BEB-B method gives good agreement with TICS computed using the BEB model. For imidazole and 2-nitroimidazole, we could not find any experimental PICS study, and hence are computed in the present study for the first time. Hence, more investigation both theoretical and experimental is needed for these molecules. The appearance energies and relative abundances of the fragments were the key parameters for the determination of the PICS, which were all taken from the experimental papers for our calculations. All other parameters such as the experimental and theoretical BRs, scaling factors, and structural factors employed for the determination of PICS were computed and are presented in Tables. 1, 2, 3 and 4 respectively. The TICS data obtained after summing all the PICS of the fragments in the MSD and m-BEB-B method underestimates the TICS data computed using the BEB model for imidazole and 2-nitroimidazole. The discrepancy is due to the lack of complete information on the fragmentation of the parent molecule, where only the major dissociation fragments are given for imidazole and 2-nitroimidazole.

### 22 Dec 2023 v2

@@ III.1 R-Carvone
In Figure 5(b), we have compared all the TICS from the existing literature along with our data. Our BEB TICS compares very well with the BEB TICS presented by Tejas et al.Jani et al. (2023), albeit their TICS calculated using the csp-ic method has a higher magnitude in comparison to BEB data, especially at higher energies after the peak. The magnitudes of the sums of the PICS calculated using both the MSD and the mBEB-B methods are lower than the TICS due to the lack of information on all the fragments and the AEs as our methods are dependent on the AEs and mass spectrum data of the fragments of the parent molecule.
@@
The PICS of 2NI is shown in Figure. 6(a), which contains the PICS of 8 fragments. The smallest fragment is of m/z of 28 which has an AE of 11.44 eV and the biggest fragment is having the m/z of 97. The cation has two AEs. Like the PICS of IMI a similar trend is seen here, both the MSD PICS and the mBEB-B PICS show a good agreement from the ionization threshold to 100 eV, after which the PICS of the MSD method falls short on comparison with the mBEB-B PICS, but their magnitudes remain similar.
@@
The TICS of 2NI is shown in Figure. 6(b), with the BEB model is compared with the TICS obtained after summing the cross sections of all the available fragments in the MSD and mBEB-B methods. The TICS do not match very well with the BEB TICS, as we don’t have information on the AEs of all the fragments that appeared in the EIMS. All the details of the parameters required to compute PICS are given in the Table. 4, which contains the theoretical BRs, the experimental BRs, the scaling factors, and the structural factors along with the maximum CS.
@@ IV Conclusion
In the present work, we have theoretically computed the partial and total ionization cross sections of R-carvone, 2-butanol, imidazole, and 2-nitroimidazole using the MSD and mBEB-B methods in combination with the BEB model. We obtained a fairly good agreement of our theoretical PICS data of various fragments with the experimental PICS for R-carvone and 2-butanol. Also, the TICS which is obtained as the sum of all the PICS of the fragments in the MSD and mBEB-B method gives good agreement with TICS computed using the BEB model. For imidazole and 2-nitroimidazole, we could not find any experimental PICS study, and hence are computed in the present study for the first time. Hence, more investigation both theoretical and experimental is needed for these molecules. The appearance energies and relative abundances of the fragments were the key parameters for the determination of the PICS, which were all taken from the experimental papers for our calculations. All other parameters such as the experimental and theoretical BRs, scaling factors, and structural factors employed for the determination of PICS were computed and are presented in Tables. 1, 2, 3 and 4 respectively. The TICS data obtained after summing all the PICS of the fragments in the MSD and mBEB-B method underestimates the TICS data computed using the BEB model for imidazole and 2-nitroimidazole. The discrepancy is due to the lack of complete information on the fragmentation of the parent molecule, where only the major dissociation fragments are given for imidazole and 2-nitroimidazole.